# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v1

@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
Our distributed-memory HPC implementation of ACCORD-FBS, termed HP-ACCORD, iteratively conducts the sparse-dense matrix multiplication for the gradient step as a special case of the SpDM3 algorithm (Koanantakool et al., 2018). This algorithm, named one-dimensionally distributed matrix multiplication (1DMM), is summarized in Algorithm 2. In HP-ACCORD, $\boldsymbol{\Omega}$ and $\boldsymbol{S}$ are separated by columns (row-wise separation is also possible for $\boldsymbol{\Omega}$) and store each partition in different computational nodes. Then, the gradient $\nabla g(\boldsymbol{\Omega})$ is computed with 1DMM. While the standard method for distributed matrix-matrix multiplication on HPC systems is the scalable universal matrix multiplication algorithm (SUMMA, van de Geijn and Watts, 1997), this method partitions involved matrices by both rows and columns and allocates them across computational nodes, and completing each block of the product requires multiple rounds of broadcasting submatrices of both operands, which results in a major performance bottleneck. On the other hand, 1DMM sends and receives one-dimensional blocks of only one operand, i.e., $\boldsymbol{\Omega}$ in Algorithm 2. Because of this difference, the communication cost of Algorithm 2 is smaller than SUMMA since $\boldsymbol{\Omega}$ is sparse while $\boldsymbol{S}$ is dense (Koanantakool et al., 2016). Computation of the gradient can alternatively be conducted in two steps: $\boldsymbol{Y}=\boldsymbol{\Omega}\boldsymbol{X}^{T}$ and $\nabla g(\boldsymbol{\Omega})=(1/n)\boldsymbol{Y}\boldsymbol{X}$, each of which can be computed with 1DMM. This strategy is advantageous when $n$ is much smaller than $p$ (Koanantakool et al., 2018).
@@ 3.4 Tuning
The choice of the regularization parameter $\lambda$ impacts the practical performance of the ACCORD estimator. At the omics scale, sample reuse methods such as cross validation or neighborhood selection (Meinshausen and Bühlmann, 2010) are ruled out since they incur a nontrivial number of expensive passes of computing the estimator. Following CONCORD and SPACE, which are also pseudolikelihood-based methods, we adopt a Bayesian information criterion (BIC)-type approach for tuning. Specifically, using the loss function part of (3.2), we choose $\lambda$ minimizing
@@
where $\|\hat{\boldsymbol{\Omega}}\|_{0}$ is the number of nonzero off-diagonal elements of the estimate $\hat{\boldsymbol{\Omega}}=\hat{\boldsymbol{\Omega}}(\lambda)$. The last term is taken from the extended BIC for graphical lasso (Foygel and Drton, 2010), proposed to promote further sparsity; $\gamma\in(0,1]$ is a user-specified parameter. We may call quantity (3.7) an extended pseudo-BIC (epBIC).
@@ 3.5 Bias correction
Correcting the biases introduced by the $\ell_{1}$ penalization by refitting is a common practice (Meinshausen, 2007; Candes and Tao, 2007; Cai, Liu and Luo, 2011). Following these approaches, we propose the following second-stage refitting procedure for ACCORD. If we let the support of the ACCORD estimate $\hat{\boldsymbol{\Omega}}$ computed by solving (3.2) for an appropriate $\lambda$ (e.g., epBIC of (3.7)) be $\hat{S}_{\lambda}$, then we refit by computing
@@ 4.1 Estimation error bounds
The consistency of $\hat{\boldsymbol{\Omega}}$ in terms of $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$ is obtained if $\sigma^{2}$ and $\kappa_{\Omega^{\ast}}$ are bounded, at the rate of $O_{P}(\sqrt{n^{-1}|S|\log p})$. The latter is bounded if, e.g., the graph implied by $\boldsymbol{\Theta}^{*}$ has a finite maximum degree of $d=\max_{i\in[p]}|\{j\in[p]:\allowbreak\theta^{\ast}_{ij}\neq 0\}|$: since $|\theta^{*}_{ij}|\leq\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$, $\kappa_{\Omega^{\ast}}=\max_{i\in[p]}\sum_{j=1}^{p}|\theta^{*}_{ij}|/\sqrt{\theta^{*}_{ii}}\leq d\max_{i\in[p]}\sqrt{\theta^{*}_{ii}}\leq d/\sqrt{\lambda_{\min}(\boldsymbol{\Sigma}^{*})}$, where $\lambda_{\min}(\boldsymbol{\Sigma}^{*})\geq\alpha$ is the minimum eigenvalue of $\boldsymbol{\Sigma}^{*}$. This rate and the sample complexity of $n\gtrsim|S|\log p$ match those for the graphical lasso (Wainwright, 2019, Proposition 11.9).
@@ 4.2 Edge selection and sign consistency
With other quantities held fixed, the sample size required to achieve the rate $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\|_{\infty}=O_{P}(\sqrt{(\tau\log p)/n})$ is $n\gtrsim d^{2}\tau\log p$. It can be also shown that for data distributions with a bounded $4m$-th moment, the error rate is $O_{P}(\sqrt{p^{\tau/m}/n})$ for $n\gtrsim d^{2}p^{\tau/m}$; see Supplementary Material A.6. These sample complexities and rates of convergence match those of graphical lasso, obtained by Ravikumar et al. (2011) under similar conditions.
@@ 5.2 Scalability of HP-ACCORD
We next investigate the scalability of HP-ACCORD (§3.3) using simulated data of dimension $p$ up to one million. At this scale, even simulating multivariate Gaussian samples becomes a nontrivial task (Vono, Dobigeon and Chainais, 2022). We took a similar approach to the numerical experiments in Li et al. (2010): generate a sparse $p\times p$ unit lower triangular matrix $\boldsymbol{L}$ taking values in $[-1,1]$ and compute $\boldsymbol{y}=\boldsymbol{L}^{-T}\boldsymbol{x}$, $\boldsymbol{x}\sim N(\mathbf{0},\boldsymbol{I}_{p})$, by backsubstitution so that $\boldsymbol{y}$ follows $N(0,(\boldsymbol{L}\boldsymbol{L}^{T})^{-1})$. The location of off-diagonal non-zero entries of $\boldsymbol{L}$ were uniformly selected so that the graph implied by the precision matrix $\boldsymbol{L}\boldsymbol{L}^{T}$ had an average degree of $10.3$ for each $p$, and the maximum degree ranged from 42 to 66.
@@
We report the results in Figure 2. Since the memory of a single KNL node (96GB) was not enough for the computation with dimensions greater than or equal to $100,000$, an appropriate number of nodes that can handle the input size data was used for HP-ACCORD, showcasing the scalability of the algorithm. The supercomputing center imposed a 48-hour restriction on the running time of a single job. Thus we only report accurate timing for the processes finished within this limit. In panel (a), the regularization parameter $\lambda$ was set large enough so that the resulting precision matrix estimate $\hat{\boldsymbol{\Theta}}$ becomes diagonal. In panel (b), $\lambda$ was adjusted so that the numbers of nonzero entries of $\hat{\boldsymbol{\Theta}}$ have a similar scale to their true precision matrices. Note that fastclime had to use the time near the limit to estimate a precision matrix to compute estimate for two $\lambda$’s for $p=10,000$, and it could not run with data that has $p=30,000$ and above due to the memory limit. Also, for the both panel, we could observe that HP-ACCORD outperforms BigQUIC in terms of computation time, even when only a single node is employed. By adopting multiple computational nodes, HP-ACCORD was able to deliver the estimate within the time budget in all scenarios. On the contrary, BigQUIC could not finish the process on time when the dimension exceeds 100,000 and the $\lambda$ is chosen so that the estimate is non-diagonal. Internally, BigQUIC tries to find a permuted partition $\{B_{1},\cdots,B_{k}\}$ of the $p$ coordinates using a graph clustering algorithm (Karypis and Kumar, 1998; Dhillon, Guan and Kulis, 2007) so that most of the coordinate updates are performed in the diagonal blocks $(B_{1},B_{1}),\cdots,(B_{k},B_{k})$. However, the size of each partition is limited due to the memory limit of a single computational node (about 20,000 in the Nurion environment in which each node has 96GB of memory), and clustering with such a limited block size bares many off-block diagonal edges in our simulated data. In consequence, the conjugate gradient method employed to compute the off-diagonal blocks are hardly skipped, causing the algorithm to run extremely slowly. Even when $\lambda$ is so large that the estimate should be diagonal, BigQUIC failed to complete the computation if the dimension reached one million.
@@ 5.3 Edge detection and estimation error
We examine the edge detection and parameter estimation performance of ACCORD on simulated data, comparing it with CONCORD, graphical lasso (GLASSO), SPACE, CLIME. Following the hub graph setting in Peng et al. (2009), we generated a graph with 10 clusters each having 100 nodes and 90 edges and connected 100 randomly chosen node pairs from adjacent clusters. The clusters had either a hub network or scale-free graph structure. A hub network was generated by (1) constructing an Erdös-Rényi graph with 97 nodes and 45 edges, and (2) consecutively adding 3 hub nodes, each connected to random 15 nodes. A scale-free graph was generated so that the degree $k$ of the nodes has a probability $P(k)\sim k^{-2.3}$. We additionally considered a simple Erdös-Rényi graph with 1,000 nodes and 1,000 edges without the clustered structure. The partial correlation matrix corresponding to the graph was chosen so that the matrix has minimum eigenvalue of at least 0.2, and each nonzero entries have absolute value of 0.1 at least. Model selection was carried out using epBIC (3.7) for ACCORD and CONCORD for an evenly spaced grid of $\lambda$’s in logarithmic scale, while extended Gaussian BIC was used for GLASSO and CLIME. Although cross validation was used to select the $\lambda$ in the original paper of CLIME (Cai, Liu and Luo, 2011), we used BIC instead to avoid the repetitive parameter fitting. For SPACE, we used “BIC-type criterion” defined in Peng et al. (2009). For CONCORD, the matrix $\hat{\boldsymbol{\Theta}}$ minimizing (2.2) was treated as a precision matrix despite the possible inconsistency (see §2). This procedure was repeated 50 times. For ACCORD, we also considered debiasing procedure (3.8).
@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
We used the LIHC dataset consisting of expression data of 15,598 protein-coding genes (mRNAs), expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands positioned in the upstream regulatory regions of genes. These are a subset of variables provided in the original data from the TCGA data portal (Ally et al., 2017), where only the DNA methylation probes in regulatory regions or gene bodies of annotated genes were included. Since DNA hypo- and hyper-methylation status tends to be correlated locally in genomic neighborhoods and those probesets for methylation status make up a majority of the variables in the data, the underlying graph is expected to reflect a sparse precision matrix with a large number of small block diagonals if the variables were ordered by chromosomes and by within-chromosomal genomic coordinates. Most causally implicated correlations between methylation probes and mRNAs are also likely to be local with the exception of DNA methyltransferase genes and the likes, and they would not alter the presumed network topology described above. The same logic also applies to mRNA-mRNA correlations, but these correlations can be local and global across genomic locations. However, these latter correlations tend to form block diagonal patterns by their upstream TF regulators and the target precision matrix will still remain within a structure amenable to consistent estimation.
@@ 6.3 Dissecting active regulatory contribution of TFs and repressive epigenomic regulation
Since the impact of methylation-driven co-repression of mRNA expression was accounted for, the adjusted network data allows us to infer active co-regulation driven by TFs and co-activators without potential confounding by epignomic regulation. We thus visualized the adjusted network using the Cytoscape software (Shannon et al., 2003) and selected six subnetworks of high connectivity (Figure 4C). We performed hypergeometric probability-based tests for enrichment of biological functions and TFs potentially implicated in each subnetwork of genes. Indeed, the subnetworks showed highly specific enrichment of widely-known cancer associated biological processes, ranging from the extracellular matrix remodeling in the parenchymal environment, immune response and signal transduction, cell cycle and protein translation, and small molecular metabolism with previously well characterized TFs in various cellular processes.
By contrast, the unadjusted network showed a higher degree of connectivity between the nodes (visualization not shown due to lack of legibility) and the overall network lacked clear separation of subnetworks. Interestingly, the network had a large number of nodes with one or two edges connecting to others only. To investigate the quality of the two networks in terms of the recovery of well-validated TF-target relationships, we queried the selected mRNA-mRNA edges in the two networks against bona fide TF-target regulatory network compiled from credible sources (Koh et al., 2019). Upon mapping the two mRNA-mRNA networks, 32% and 30% of the edges involved genes encoding a TF or a co-activator in the adjusted and unadjusted networks, respectively. However, while 19% of the reported edges were experimentally validated TF-target pairs in the adjusted network, only 9% of the edges were validated in the unadjusted network. Given the unadjusted network is about 2.6 times larger than the adjusted network, the result implies that the adjusted network recovers the known TF-target regulatory relationships with comparable sensitivity and greater specificity.

### v2

@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
Our distributed-memory HPC implementation of ACCORD-FBS, termed HP-ACCORD, iteratively conducts the sparse-dense matrix multiplication for the gradient step as a special case of the SpDM3 algorithm [20]. This algorithm, named one-dimensionally distributed matrix multiplication (1DMM), is summarized in Algorithm 2. In HP-ACCORD, $\boldsymbol{\Omega}$ and $\boldsymbol{S}$ are separated by columns (row-wise separation is also possible for $\boldsymbol{\Omega}$) and store each partition in different computational nodes. Then, the gradient $\nabla g(\boldsymbol{\Omega})$ is computed with 1DMM. While the standard method for distributed matrix-matrix multiplication on HPC systems is the scalable universal matrix multiplication algorithm [SUMMA, 36], this method partitions involved matrices by both rows and columns and allocates them across computational nodes, and completing each block of the product requires multiple rounds of broadcasting submatrices of both operands, which results in a major performance bottleneck. On the other hand, 1DMM sends and receives one-dimensional blocks of only one operand, i.e., $\boldsymbol{\Omega}$ in Algorithm 2. Because of this difference, the communication cost of Algorithm 2 is smaller than SUMMA since $\boldsymbol{\Omega}$ is sparse while $\boldsymbol{S}$ is dense [19]. Computation of the gradient can alternatively be conducted in two steps: $\boldsymbol{Y}=\boldsymbol{\Omega}\boldsymbol{X}^{T}$ and $\nabla g(\boldsymbol{\Omega})=(1/n)\boldsymbol{Y}\boldsymbol{X}$, each of which can be computed with 1DMM. This strategy is advantageous when $n$ is much smaller than $p$ [20].
@@ 3.4 Tuning
The choice of the regularization parameter $\lambda$ impacts the practical performance of the ACCORD estimator. At the omics scale, sample reuse methods such as cross validation or neighborhood selection [28] are ruled out since they incur a nontrivial number of expensive passes of computing the estimator. Following CONCORD and SPACE, which are also pseudolikelihood-based methods, we adopt a Bayesian information criterion (BIC)-type approach for tuning. Specifically, using the loss function part of (3.2), we choose $\lambda$ minimizing
@@
where $\|\hat{\boldsymbol{\Omega}}\|_{0}$ is the number of nonzero off-diagonal elements of the estimate $\hat{\boldsymbol{\Omega}}=\hat{\boldsymbol{\Omega}}(\lambda)$. The last term is taken from the extended BIC for graphical lasso [10], proposed to promote further sparsity; $\gamma\in(0,1]$ is a user-specified parameter. We may call quantity (3.7) an extended pseudo-BIC (epBIC).
@@ 3.5 Bias correction
Correcting the biases introduced by the $\ell_{1}$ penalization by refitting is a common practice [26, 5, 4]. Following these approaches, we propose the following second-stage refitting procedure for ACCORD. If we let the support of the ACCORD estimate $\hat{\boldsymbol{\Omega}}$ computed by solving (3.2) for an appropriate $\lambda$ (e.g., epBIC of (3.7)) be $\hat{S}_{\lambda}$, then we refit by computing
@@ 4.1 Estimation error bounds
The consistency of $\hat{\boldsymbol{\Omega}}$ in terms of $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$ is obtained if $\sigma^{2}$ and $\kappa_{\Omega^{\ast}}$ are bounded, at the rate of $O_{P}(\sqrt{n^{-1}|S|\log p})$. The latter is bounded if, e.g., the graph implied by $\boldsymbol{\Theta}^{*}$ has a finite maximum degree of $d=\max_{i\in[p]}|\{j\in[p]:\allowbreak\theta^{\ast}_{ij}\neq 0\}|$: since $|\theta^{*}_{ij}|\leq\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$, $\kappa_{\Omega^{\ast}}=\max_{i\in[p]}\sum_{j=1}^{p}|\theta^{*}_{ij}|/\sqrt{\theta^{*}_{ii}}\leq d\max_{i\in[p]}\sqrt{\theta^{*}_{ii}}\leq d/\sqrt{\lambda_{\min}(\boldsymbol{\Sigma}^{*})}$, where $\lambda_{\min}(\boldsymbol{\Sigma}^{*})\geq\alpha$ is the minimum eigenvalue of $\boldsymbol{\Sigma}^{*}$. This rate and the sample complexity of $n\gtrsim|S|\log p$ match those for the graphical lasso [39, Proposition 11.9].
@@ 4.2 Edge selection and sign consistency
With other quantities held fixed, the sample size required to achieve the rate $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\|_{\infty}=O_{P}(\sqrt{(\tau\log p)/n})$ is $n\gtrsim d^{2}\tau\log p$. It can be also shown that for data distributions with a bounded $4m$-th moment, the error rate is $O_{P}(\sqrt{p^{\tau/m}/n})$ for $n\gtrsim d^{2}p^{\tau/m}$; see Supplementary Material A.6. These sample complexities and rates of convergence match those of graphical lasso, obtained by Ravikumar et al. [33] under similar conditions.
@@ 5.2 Scalability of HP-ACCORD
We next investigate the scalability of HP-ACCORD (§3.3) using simulated data of dimension $p$ up to one million. At this scale, even simulating multivariate Gaussian samples becomes a nontrivial task [38]. We took a similar approach to the numerical experiments in Li et al. [23]: generate a sparse $p\times p$ unit lower triangular matrix $\boldsymbol{L}$ taking values in $[-1,1]$ and compute $\boldsymbol{y}=\boldsymbol{L}^{-T}\boldsymbol{x}$, $\boldsymbol{x}\sim N(\mathbf{0},\boldsymbol{I}_{p})$, by backsubstitution so that $\boldsymbol{y}$ follows $N(0,(\boldsymbol{L}\boldsymbol{L}^{T})^{-1})$. The location of off-diagonal non-zero entries of $\boldsymbol{L}$ were uniformly selected so that the graph implied by the precision matrix $\boldsymbol{L}\boldsymbol{L}^{T}$ had an average degree of $10.3$ for each $p$, and the maximum degree ranged from 42 to 66.
@@
We report the results in Figure 2. Since the memory of a single KNL node (96GB) was not enough for the computation with dimensions greater than or equal to $100,000$, an appropriate number of nodes that can handle the input size data was used for HP-ACCORD, showcasing the scalability of the algorithm. The supercomputing center imposed a 48-hour restriction on the running time of a single job. Thus we only report accurate timing for the processes finished within this limit. In panel (a), the regularization parameter $\lambda$ was set large enough so that the resulting precision matrix estimate $\hat{\boldsymbol{\Theta}}$ becomes diagonal. In panel (b), $\lambda$ was adjusted so that the numbers of nonzero entries of $\hat{\boldsymbol{\Theta}}$ have a similar scale to their true precision matrices. Note that fastclime had to use the time near the limit to estimate a precision matrix to compute estimate for two $\lambda$’s for $p=10,000$, and it could not run with data that has $p=30,000$ and above due to the memory limit. Also, for the both panel, we could observe that HP-ACCORD outperforms BigQUIC in terms of computation time, even when only a single node is employed. By adopting multiple computational nodes, HP-ACCORD was able to deliver the estimate within the time budget in all scenarios. On the contrary, BigQUIC could not finish the process on time when the dimension exceeds 100,000 and the $\lambda$ is chosen so that the estimate is non-diagonal. Internally, BigQUIC tries to find a permuted partition $\{B_{1},\cdots,B_{k}\}$ of the $p$ coordinates using a graph clustering algorithm [17, 9] so that most of the coordinate updates are performed in the diagonal blocks $(B_{1},B_{1}),\cdots,(B_{k},B_{k})$. However, the size of each partition is limited due to the memory limit of a single computational node (about 20,000 in the Nurion environment in which each node has 96GB of memory), and clustering with such a limited block size bares many off-block diagonal edges in our simulated data. In consequence, the conjugate gradient method employed to compute the off-diagonal blocks are hardly skipped, causing the algorithm to run extremely slowly. Even when $\lambda$ is so large that the estimate should be diagonal, BigQUIC failed to complete the computation if the dimension reached one million.
@@ 5.3 Edge detection and estimation error
We examine the edge detection and parameter estimation performance of ACCORD on simulated data, comparing it with CONCORD, graphical lasso (GLASSO), SPACE, CLIME. Following the hub graph setting in Peng et al. [32], we generated a graph with 10 clusters each having 100 nodes and 90 edges and connected 100 randomly chosen node pairs from adjacent clusters. The clusters had either a hub network or scale-free graph structure. A hub network was generated by (1) constructing an Erdös-Rényi graph with 97 nodes and 45 edges, and (2) consecutively adding 3 hub nodes, each connected to random 15 nodes. A scale-free graph was generated so that the degree $k$ of the nodes has a probability $P(k)\sim k^{-2.3}$. We additionally considered a simple Erdös-Rényi graph with 1,000 nodes and 1,000 edges without the clustered structure. The partial correlation matrix corresponding to the graph was chosen so that the matrix has minimum eigenvalue of at least 0.2, and each nonzero entries have absolute value of 0.1 at least. Model selection was carried out using epBIC (3.7) for ACCORD and CONCORD for an evenly spaced grid of $\lambda$’s in logarithmic scale, while extended Gaussian BIC was used for GLASSO and CLIME. Although cross validation was used to select the $\lambda$ in the original paper of CLIME [4], we used BIC instead to avoid the repetitive parameter fitting. For SPACE, we used “BIC-type criterion” defined in Peng et al. [32]. For CONCORD, the matrix $\hat{\boldsymbol{\Theta}}$ minimizing (2.2) was treated as a precision matrix despite the possible inconsistency (see §2). This procedure was repeated 50 times. For ACCORD, we also considered debiasing procedure (3.8).
@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
We used the LIHC dataset consisting of expression data of 15,598 protein-coding genes (mRNAs), expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands positioned in the upstream regulatory regions of genes. These are a subset of variables provided in the original data from the TCGA data portal [1], where only the DNA methylation probes in regulatory regions or gene bodies of annotated genes were included. Since DNA hypo- and hyper-methylation status tends to be correlated locally in genomic neighborhoods and those probesets for methylation status make up a majority of the variables in the data, the underlying graph is expected to reflect a sparse precision matrix with a large number of small block diagonals if the variables were ordered by chromosomes and by within-chromosomal genomic coordinates. Most causally implicated correlations between methylation probes and mRNAs are also likely to be local with the exception of DNA methyltransferase genes and the likes, and they would not alter the presumed network topology described above. The same logic also applies to mRNA-mRNA correlations, but these correlations can be local and global across genomic locations. However, these latter correlations tend to form block diagonal patterns by their upstream TF regulators and the target precision matrix will still remain within a structure amenable to consistent estimation.
@@ 6.3 Dissecting active regulatory contribution of TFs and repressive epigenomic regulation
Since the impact of methylation-driven co-repression of mRNA expression was accounted for, the adjusted network data allows us to infer active co-regulation driven by TFs and co-activators without potential confounding by epignomic regulation. We thus visualized the adjusted network using the Cytoscape software [34] and selected six subnetworks of high connectivity (Figure 4C). We performed hypergeometric probability-based tests for enrichment of biological functions and TFs potentially implicated in each subnetwork of genes. Indeed, the subnetworks showed highly specific enrichment of widely-known cancer associated biological processes, ranging from the extracellular matrix remodeling in the parenchymal environment, immune response and signal transduction, cell cycle and protein translation, and small molecular metabolism with previously well characterized TFs in various cellular processes.
By contrast, the unadjusted network showed a higher degree of connectivity between the nodes (visualization not shown due to lack of legibility) and the overall network lacked clear separation of subnetworks. Interestingly, the network had a large number of nodes with one or two edges connecting to others only. To investigate the quality of the two networks in terms of the recovery of well-validated TF-target relationships, we queried the selected mRNA-mRNA edges in the two networks against bona fide TF-target regulatory network compiled from credible sources [21]. Upon mapping the two mRNA-mRNA networks, 32% and 30% of the edges involved genes encoding a TF or a co-activator in the adjusted and unadjusted networks, respectively. However, while 19% of the reported edges were experimentally validated TF-target pairs in the adjusted network, only 9% of the edges were validated in the unadjusted network. Given the unadjusted network is about 2.6 times larger than the adjusted network, the result implies that the adjusted network recovers the known TF-target regulatory relationships with comparable sensitivity and greater specificity.